Jan. 21-24

– Four stores provided ~56.75 hours of luminosity
– An LCW leak on a Pbar Accumulator magnet required an access to repair
– Shot setup held up due to noisy readbacks
– Tevatron quench during scrape in shot setup
– Linac personel switched from I Source to H Source yesterday

The integrated luminosity for the period from 1/17/11 to 1/24/11 was 70.18 inverse pictobarns. NuMi reported receiving 8.61E18 protons on target during this same period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts